OpenStudy (anonymous):

What are the different methods of writing a vector's angle/direction (eg. navigator method, etc.)?

OpenStudy (osanseviero):

One is writing in forms of the vectors i and j i=\[\left(\begin{matrix}1 \\ 0\end{matrix}\right)\]

OpenStudy (osanseviero):

j=\[\left(\begin{matrix}0 \\ 1\end{matrix}\right)\]

OpenStudy (osanseviero):

So a vector V=3i+8j is the same than \[\left(\begin{matrix}3 \\ 8\end{matrix}\right)\]

OpenStudy (osanseviero):

And those are the two forms of writing that I know

OpenStudy (kainui):

You can write a vector like this, <a,b> which means a is the x-component and b is the y-component. But if you want vectors with angles involved, then what you do is find out the length of the vector by looking at it as the hypotenuse of a right triangle.  So you can see that the vector's length is r=sqrt(x^2+y^2) and then you can also see that the angle of the vector is \[\theta = \tan^{-1}(\frac{ y }{ x})\] And don't forget x=r*cos(theta) and y=r*sin(theta) if you need to find one thing or another, they can be useful. So now you should have everything you need to represent a vector in terms of its magnitude and direction (vector length and angle).